OpenStudy (anonymous):

Type the ordered pair that is the solution to these equations. x = 5y 2x - 3y = 14

OpenStudy (callisto):

x = 5y -(1) 2x - 3y = 14 -(2) Sub. (1) into (2) 2(5y) - 3y = 14 Solve y. Then, put the value of y into (1) to get x.